OpenStudy (nikki_skellington):

A cylinder and a cone have the same diameter: 8 inches. The height of the cylinder is 3 inches. The height of the cone is 18 inches. Use π = 3.14. What is the relationship between the volume of this cylinder and this cone? Explain your answer by determining the volume of each and comparing them. Show all your work.

OpenStudy (nikki_skellington):

@Vocaloid?

Vocaloid (vocaloid):

start by finding the radius of the cone and cylinder radius = (1/2) * diameter = ?

OpenStudy (nikki_skellington):

it would be 4

Vocaloid (vocaloid):

good, now let's find the volume of the cone

Vocaloid (vocaloid):

volume = (1/3) * pi * radius^2 * height = ?

OpenStudy (nikki_skellington):

i got 96 pi

Vocaloid (vocaloid):

good now let's find the volume of the cylinder

Vocaloid (vocaloid):

volume = pi * radius^2 * height = ?

OpenStudy (nikki_skellington):

48pi?

Vocaloid (vocaloid):

good! so if the volume of the cone is 96pi and the volume of the cylinder is 48 pi that means the volume of the cone is 2 times the volume of the cylinder and that's your answer !

OpenStudy (nikki_skellington):

thank you so much!

Vocaloid (vocaloid):

just as a side note, they want you to use 3.14 for pi

Vocaloid (vocaloid):

so you would just write the volumes as 96 * 3.14 = 301.44 inches cubed for the cone and 150.72 for the cylinder

Vocaloid (vocaloid):

everything else we said before is the same though

OpenStudy (nikki_skellington):

ok thanks.